the perimeter of a rectangular garden is 4o ft if the length is 6 ft more than the width what are the length and the width of the garden

Answer by cleomenius(959)

x = length
x + 6 = height.
2x + 2x + 12 = 40
4x = 28
x = 7
13 ft long
7 ft wide.
26 + 14 = 40 feet.
Cleomenius.